OpenStudy (mathmath333):

is zero \(\huge 0\) a positive multiple of every integer ?

OpenStudy (kainui):

If by positive you mean not less than 0 and by multiple you mean a*n is a multiple of n, then yes.

OpenStudy (mathmath333):

so it is positive multiple of every integer .?

OpenStudy (mathmath333):

i m very confused

OpenStudy (kainui):

It depends on what you define positive to mean. Does positive mean greater than 0 or greater than and equal to zero? Unless your class has a specific definition of positive, this is just a vague question.

OpenStudy (mathmath333):

i dont know what exactly positive means \(0>\) or \(0\geq \) , i want an undisputed universally accepted fact.

ganeshie8 (ganeshie8):

0 is not positive, not negative

OpenStudy (mathmath333):

so that implies its neither a positive multiple ?

ganeshie8 (ganeshie8):

nonnegative integers include 0,1,2,3,... nonpositive integers include 0,-1,-2,-3,...

ganeshie8 (ganeshie8):

for quick checks, we can rely on wolfram, it has pretty consistent definitions for most number theory terms http://mathworld.wolfram.com/NonpositiveInteger.html

OpenStudy (mathmath333):

if the question asks what are the first 5 multiples of 5 ? and what are the first 5 positive multiples of 5 ? what should i write

ganeshie8 (ganeshie8):

5, 10, 15, 25, 30 are the first five positive multiples of 5

ganeshie8 (ganeshie8):

*missed 20

OpenStudy (mathmath333):

and ehat about this \(\text{what are the first 5} \cancel {positive } \text{multiples of 5 ?}\)

ganeshie8 (ganeshie8):

the question just wants you list the familiar multiplication table : \[5\times 1 = 5\\5\times2=10\\5\times 3 = 15\\5\times 4=20\\5\times 5=25\]

OpenStudy (mathmath333):

so if for the question \(\text{what are the first 5 multiples of 5}\) i should not include \(0\)

ganeshie8 (ganeshie8):

\[5\times 0 = 0\] Oh you mean you want to include that ?

OpenStudy (mathmath333):

yes that one!

ganeshie8 (ganeshie8):

yeah you may include that but just keep in mind 0 is not positive

OpenStudy (mathmath333):

ohk thnks so first five multiples of 5 are 0 ,5 ,10, 15, 20

OpenStudy (kainui):

Yeah I guess I was thinking for some reason of how there's the confusion between is 0 a natural number or not, which is a slightly different question.

ganeshie8 (ganeshie8):

i would stick to 5,10,15,20,25

ganeshie8 (ganeshie8):

thats right because \[0 = x\times 0 \] for any integer \(x\)

OpenStudy (mathmath333):

in many books, i found that they dont include 0

OpenStudy (kainui):

Yeah there are two parts going on here. You are looking for positive multiples, but if you want to look at strictly multiples of 5, this is your infinite list: \[\Large ... ,-25, -5, 0, 5, 25, ...\]

OpenStudy (mathmath333):

ok so postive multiple \(\cancel 0\) and only for only multiple \(0~~\checkmark\) i should stick to this

ganeshie8 (ganeshie8):

i guess question should clarify what it is talking about. Knowing 0 is neither positive nor negative is sufficient and rest should become clear based on context of the quesiton..

OpenStudy (mathmath333):

thnks both for clarifying

ganeshie8 (ganeshie8):

wolfram doesn't like 0 http://www.wolframalpha.com/input/?i=first+five+multiples+of+5

OpenStudy (mathmath333):

oh my god

ganeshie8 (ganeshie8):

"what are the first five multiples of 5 ?" is somewhat incomplete question... because there is no least element in the set of multiples of 5

ganeshie8 (ganeshie8):

"what are the first five nonegative multiples of 5 ?" makes sense because this set starts with 0 : {0, 5, 10, 15, ... }

OpenStudy (mathmath333):

wolfram was interpreting question as 5 between 5 and 25

ganeshie8 (ganeshie8):

lol yeah it seems mother wolfram has its own way of interpreting what you're thinking..

OpenStudy (mathmath333):

i should write an email to wolfram as perl does sometimes .

OpenStudy (mathmath333):

http://www.wolframalpha.com/input/?i=do+u+love+me+%3F

OpenStudy (mathmath333):

got the email lol http://www.wolframalpha.com/input/?i=whats+ur+email+address

ganeshie8 (ganeshie8):

its pretty amazing how it dynamically interprets the queries and responds to some of the weird questions lol xD

OpenStudy (kainui):

This is unfortunate... =( http://www.wolframalpha.com/input/?i=6accdae13eff7i3l9n4o4qrr4s8t12ux

OpenStudy (mathmath333):

what was that

OpenStudy (kainui):

It's a literal quote from Isaac Newton ;P

OpenStudy (mathmath333):

haha

OpenStudy (mathmath333):

this also doesnt include 0 http://www.wolframalpha.com/input/?i=what+are+multiples+of+5+%3F